Find an equation of the line having the given slope and containing the given point.
m= −2​, ​(6​,9​)

Answer by Alan3354(69443)

Find an equation of the line having the given slope and containing the given point.
m = -2, (6,9)
-------
y-y1 = m*(x-x1)
Just write it down:
--
y-9 = -2*(x-6)